OpenStudy (anonymous):

explain why 3x^2 + kx - 1 is never always positive for any value of k

OpenStudy (anonymous):

the formula of discriminant is b^2 - 4ac in this case: a=3 b=k c= -1 so, k^2 always has positive value, then k^2 (positive number) -4 * 3 * (-1)= k -(-12)= k^2 + 12. positive number + positive number = always positive number

OpenStudy (anonymous):

so 3x^2 + kx - 1 is always positive for any value of k